Find the solution to the following system of linear equations:
3x+y=8
-2x+y=-7

Answer by lawdej(58)

use substitution method or elimination method
for substitution method take one of the equations
-2x+y=-7
y=2x-7 substitute this into 3x+y=8
3x+2x-7=8
5x-7=8
5x=15
x=3
so y=2x-7 =2x3-7=6-7=-1
thus x=3 and y=-1